OpenStudy (anonymous):

calculate (f * g)' (3) f'(3)=-1 g'(3)=9

OpenStudy (anonymous):

Let f(x) = (3x^3)/((4-5x)^5) find the equation of the line tangent to the graph of f at x=2

OpenStudy (anonymous):

what question is tht for what section?

OpenStudy (anonymous):

chain rule problem 8

OpenStudy (anonymous):

haven't done tht one yet

OpenStudy (anonymous):

it's due now

OpenStudy (anonymous):

yeah i know

myininaya (myininaya):

(fg)'=f'g+fg' (fg)'(3)=f'(3)g(3)+f(3)g'(3) =-g(3)+9f(3) anything else given is that f times g or f(g(x))

myininaya (myininaya):

i assume it was times

OpenStudy (anonymous):

yea it was times the f and g

OpenStudy (anonymous):

thanks for the step by step too!!